OpenStudy (anonymous):

If m =3 + 2p^2 and n=4 + p(sqrt2), what is m in terms of n? A. 4n^2 + 16n + 32 B. 4n^2 + 16n(sqrt2) + 29 C. 4n^2 + 16n(sqrt2) + 35 D. n^2 - 8n +19 E. n^2 +8n +19 Can someone help me figure out how to set up for this problem? I don't know where to start.

Nnesha (nnesha):

you have to solve second equation for p

Nnesha (nnesha):

\[n = 4 + p \sqrt{2} \] second equation this right ??

Nnesha (nnesha):

if your equation look like this so subtract 4 both side and then divide sqrt 2

OpenStudy (anonymous):

So \[n-4/\sqrt{2}=p\]?

Nnesha (nnesha):

yes now plug into the first equation where p is

OpenStudy (anonymous):

m = 3 + 2(n-4/(sqrt2)^2

Nnesha (nnesha):

yes good so so you can write that like this